The odometer of a car reads 57321.0km when the clock shows the time 08:30 AM. What is the distance moved by the car, if at 08:50AM, the odometer reading has changed to 57336.0km?

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this class X maths answer. We will update the answer very soon.
Yash Patari
2 years ago
DEAR STUDENT
Therefore, Distance travelled = (57336.0-57321.0) km = 15.0 km, Time taken = 8:50 am – 8:30 am = 20 min. = 0.75 × 60 km/h = 45 km/h.
THANKS